# Pay Attention to Your Neighbours: Training-Free Open-Vocabulary Semantic Segmentation

### v1

@@ 3.2 Limitations of CLIP for Image Segmentation
As previously stated, localization holds a pivotal role in segmentation tasks, yet it is often overlooked in ViTs [10]. As described in [10], unlike CNNs, most of the operations in ViTs are global and the locality of patches is mostly not taken into account. Within ViT, and consequently, in CLIP’s visual encoder, positional information is integrated into the network through 1D learnable positional embeddings added to the input. We contend that three issues arise with this approach, particularly in dense prediction tasks. Firstly, when dealing with images, we encounter two dimensions, making it suboptimal to rely solely on a 1D positional embedding. Secondly, unlike the original transformer [31], which employs $\sin$ and $\cos$ functions to compute positional embeddings (enabling it to encapsulate information about relative distances), ViT utilizes randomly initialized learnable parameters. Although these parameters could theoretically learn to prioritize attention between adjacent patches, there is no guarantee that they will indeed do so. While this might not pose a significant issue in classification, where only the representation of the [CLS] token matters, attending to adjacent patches in semantic segmentation becomes crucial, as for the vast majority of patches they contain important contextual information. Lastly, even if positional embeddings encode information about patch locality, their utilization is restricted to the first encoder block, potentially diminishing their relevance in subsequent layers. This oversight could be detrimental in dense prediction tasks where such information remains valuable throughout the network’s depth. In light of the arguments presented above, we assert that explicitly attending to the neighbourhood of each patch is imperative within the context of segmentation. Moreover, recent studies have noted that the final encoder block in CLIP’s ViT disrupts spatial information, hindering dense prediction tasks [46]. Indeed, CLIP’s visual encoder is trained to emphasize the [CLS] token’s embedding (global embedding), while the outputs at other locations (i.e., embeddings of patches) are not optimally structured for tasks such as semantic segmentation.
@@ 4 Neighbour-Aware CLIP
The pre-training procedure of CLIP [27] encourages its vision transformer [10] to learn representations tailored to image-level tasks, thereby compromising its efficacy in dense prediction problems. Given the inherent differences between such tasks and segmentation, coupled with the fact that the output of the ViT for patch tokens has not been explicitly trained during CLIP’s pre-training, it struggles to generalize effectively to pixel-wise prediction scenarios. This underscores the necessity for targeted adjustments in the original CLIP model to accommodate the nuances of semantic segmentation. Our study delves into these specific components that potentially hinder CLIP’s segmentation performance and proposes minimal alterations to its overall framework, without changing any of the network’s parameters. The precise modifications are detailed in this section.
Removing the [CLS] token.
Before discussing the proposed changes, for the sake of notation simplicity, we consider removing the [CLS] token. Although it plays a primary role in the original CLIP’s encoders, our efforts to extract from it information useful for segmentation proved unsuccessful. We contend that this failure can be attributed to the divergent nature of CLIP’s pre-training and the specific requirements of dense prediction problems, in which objects of interest have to be recognized with their local positioning. Given that we do not use this token in our approach, eliminating it leads to more straightforward formulations. Without [CLS], the following statements hold about Eqs. 5, 6, 7 and 8:
$\mathbf{Z}$ $\in{\mathbb{R}}^{h\times w\times D},$ $\mathbf{W}^{qkv}$ $\in{\mathbb{R}}^{D\times d\times 3},$ $\mathbf{q},\,\mathbf{k},\,\mathbf{v}$ $\in{\mathbb{R}}^{h\times w\times d},$ (9)
$\text{sim}_{ij}$ $\in{\mathbb{R}}^{h\times w},$ $\mathbf{W}^{o}$ $\in{\mathbb{R}}^{d\times D},$ $\operatorname{SA}(\mathbf{Z})$ $\in{\mathbb{R}}^{h\times w\times D}.$ (10)
@@ 4.1 Introducing Spatial Consistency
In Sec. 3.2, we underscored the importance of explicitly attending to the locality of each patch and highlighted the inadequacy of vanilla ViT’s positional embedding in achieving this. In this section, we introduce our simple method for enforcing explicit spatial attention to each patch’s neighbourhood. In particular, we augment the attention map information with an unnormalized multivariate (in our case, $2$-dimensional) Gaussian kernel, which can be defined as:
@@
If we assume $\boldsymbol{\Sigma}=\sigma^{2}I$, we can rewrite the kernel as
@@
$\begin{gathered}\omega((i,j);\sigma)_{mn}=\phi((m,n);{(i,j)},\sigma),\\ \phantom{.}\forall m\in\{1,2,\dots,h\},\;\forall n\in\{1,2,\dots,w\}.\end{gathered}$ (13)
As a corner case, let us assume that the attention map’s logits for a given patch have been explicitly set to 00 and we set
@@
so that the attention will be purely paid to the neighbourhood of patches. As we show in our empirical validation (termed Neighbourhood Only in Sec. 5.3.1), by just doing so we can observe a leap in performance compared to CLIP.
[…]
@@ 4.2 Measure of Similarity
Transformers [31] are complex deep neural networks and providing a definitive explanation of their inner workings remains challenging. However, from an intuitive standpoint, we can offer the following interpretations for query, key, and value vectors. The query vector denotes what a patch looks for; the key vector signifies what it represents; and the value vector displays what it has to offer. Guided by these descriptions, we deviate from the standard similarity measure ($\mathbf{q}\mathbf{k}^{\top}$) in our self-attention module. This deviation is motivated by the discrepancy between what we want the model to look for (i.e., accurate patch-level predictions) and what it had been trained to look for during pre-training. In semantic segmentation, we need to focus on the nature of each patch and this naturally leads us to shift our attention toward using the key vectors. Thus, we opted to use $\mathbf{k}\mathbf{k}^{\top}$ scores in our similarity measure instead, resulting in
$\phantom{,}\text{sim}_{ij}=\dfrac{\mathbf{k}\mathbf{k}_{ij}}{\sqrt{d}},$ (17)
in which $1/\sqrt{d}$ is the same scaling factor used in CLIP, as in Eq. 6. By doing so, patches that represent similar information (as portrayed by their key vectors) attend to each other’s value vectors with high intensity.
4.3 Eliminating Image-Level Specialized Components
@@
in which $L$ denotes the index of the last encoder block. We refer to this, as the Reduced architecture of the last encoder block.
@@ 5 Experiments
5.1 Experimental Setup
5.1.1 Datasets.
We evaluate our method on the following segmentation benchmarks, whose names are abbreviated (in parentheses) to conserve table space: PASCAL VOC 2012 (V21) [11], ADE20K-150 (ADE) [45], PASCAL Context (PC60) [25], COCO-Stuff (C-Stf) [5], Cityscapes (City) [9], COCO-Object (C-Obj) [20]. Additionally, alongside the original benchmarks on these datasets, we follow [32] and evaluate on variants of PASCAL VOC 2012 (V20) and PASCAL Context (PC59) in which the background class is removed from the evaluation. Furthermore, input images are resized to have a shorter side of 336 (560 for Cityscapes [9], because of high-resolution images), and a slide inference is carried out with a $224\times 224$ window and a stride of 112, following the procedure outlined in [32].
5.1.2 Baselines.
We compare our method to a set of relevant works in OVSS, including: MaskCLIP [46], ReCo [28], CLIP Surgery [18], SCLIP, [32], GEM [4], CLIP-DIY [35], and FOSSIL [3]. We also include a few influential weakly supervised OVSS methods, such as GroupViT [36] and TCL [7] in our comparison. Furthermore, since vanilla CLIP [27] can be adapted for semantic segmentation, we include it as a baseline in our comparison tables as well.
5.1.3 Implementation Details.
In our experiments, we employ pre-trained CLIP-ViT [27]. Unless mentioned otherwise, we use the ViT-B/16 backbone ($16\times 16$ patch size) comprising 12 visual encoder blocks. Since our approach operates in a training-free manner, we exclusively utilize the frozen CLIP model without any optimization. The standard deviation of the Gaussian kernel, $\sigma$ in Eq. 12, is set to 5, a choice further elaborated upon in Appendix 0.B. OVSS methods often incorporate a mask refinement step [3, 32, 7, 28], such as DenseCRF [17] or pixel-adaptive mask refinement (PAMR) [1]. For our method, we opt for PAMR as it is lighter and more efficient. We also report results without this refinement step. We resort to mIoU as the evaluation metric across all experiments.
5.2 Main Results
5.2.1 Comparison to Training-Free OVSS Methods.
In Tab. 1, we evaluate our proposed method against OVSS baselines mentioned in Sec. 5.1. Although most methods are presented as training-free OVSS, it is important to note that they vary in underlying characteristics. Specifically, within the subset of methods that refrain from fine-tuning, some leverage auxiliary pre-trained models [28, 3, 35]. Given our primary focus on achieving training-free OVSS exclusively through a frozen CLIP model, comparing our method to those utilizing additional knowledge or data, while informative, may not be entirely equitable. Hence, we explicitly denote fair comparisons in Tab. 1. Analyzing the results, our approach outperforms state-of-the-art training-free OVSS methods in 7 out of 8 benchmarks. This signifies the effectiveness of our proposed architecture and formulation of CLIP’s visual encoder. Note that results for newer methods [3, 35, 4, 18] are extracted from their respective manuscripts, hence might not encompass experiments across all benchmarks, whereas other results have been derived from [32].
Table 1: Quantitative evaluation on 6 datasets and 2 variants. The first 3 benchmarks (V21, PC60, and C-Obj) contain a background category, while the subsequent ones do not. NACLIP represents our method. The Fair column indicates whether a method’s comparison to ours is equitable, i.e., conducted without leveraging additional knowledge. For methods employing a post-processing mechanism, such as DenseCRF [17] or PAMR [1], we have included their enhanced results after mask refinement. For elucidation on abbreviated dataset names, please refer to Sec. 5.1.
Method Fair V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg
CLIP [27] ICML’21 18.6 7.8 6.5 49.1 6.7 11.2 3.2 7.2 13.8
@@ 5.2 Main Results
CLIP Surgery [18] Arxiv’23 41.2 30.5 - - 31.4 - 12.9 21.9 -
SCLIP [32] Arxiv’23 61.7 31.5 32.1 83.5 34.1 36.1 17.8 23.9 40.1
GEM [4] CVPR’24 46.2 - - - - 32.6 15.7 - -
CLIP-DIY [35] WACV’24 59.0 - 30.4 - - - - - -
@@
NACLIP [PAMR] Ours 58.8 32.2 33.2 77.1 35.5 35.2 17.4 22.9 39.0
NACLIP Ours 62.4 35.0 36.2 80.6 38.3 38.4 19.1 25.2 41.9
5.2.2 Robustness to Visual Backbones.
Table 2 reports the segmentation results of NACLIP using different CLIP-ViT backbones, namely ViT-B/16 (default), ViT-B/32, and ViT-L/14. Compared to concurrent approaches [32, 4], our method is more robust to the backbone choice. For example, SCLIP [32] experiences a significant performance drop of over 12% on average when using ViT-L/14 instead of the default backbone, whereas the degradation is less than 3% in our case. Moreover, our method outperforms GEM [4] by over 10% across all settings, making NACLIP a more robust solution compared to [32, 4].
@@
NACLIP(Ours) 62.9 38.5 50.7 57.7 34.9 46.3 59.3 36.4 47.9
5.2.3 Visual Examples.
Figure 2 presents several segmentation maps generated by our method on the PASCAL Context (59) dataset [25]. These visualizations demonstrate NACLIP’s significant performance enhancement over CLIP. Furthermore, SCLIP seems to find difficulties in identifying the limits of objects properly, sometimes confusing nearby concepts (e.g., the first row). Besides, while SCLIP tends to focus mainly on similar patches without explicitly considering the surrounding context, our method maintains a local contextual understanding. This is evident in the second row (the bus image) where SCLIP misclassifies trees, likely due to its limited attention to nearby objects such as the road.
Figure 2: Qualitative results on PASCAL Context (59) [25]. Example segmentation maps of vanilla CLIP [27], SCLIP [32], and our method are presented.
@@ 5.3 Ablation Study
5.3.1 Effect of Altering the Self-Attention Module.
In Sec. 4.1, we discussed the rationale behind introducing spatial consistency to attention maps and described our approach. Besides, we defined the Neighbourhood Only (N-Only) setting, where patches exclusively attend to their neighbours, disregarding similarities. We introduce an additional setting termed Gaussian Augmented Vanilla (GAV), wherein the Gaussian window $\omega$ is added to the similarity measure of vanilla CLIP, i.e., $\mathbf{q}\mathbf{k}^{\top}$. Results from the comparative analysis (Tab. 3) between N-Only and Vanilla, as well as between $\widetilde{A}$ (Eq. 15) and GAV, underscore the efficacy of the modifications proposed in Secs. 4.1 and 4.2, respectively.
Table 3: Ablation on the effect of spatial consistency and similarity measure. Vanilla refers to CLIP’s self-attention module and the rest of the settings have been defined in Sec. 5.3.1. Throughout this experiment, we adhere to the default CLIP encoder architecture, retaining all architectural elements of the encoder block. To provide clarity on the effects, we present results both with and without mask refinement.
Attention PAMR V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg Vanilla 18.6 7.8 6.5 49.1 6.7 11.2 3.2 7.2 13.8 N-Only 23.0 11.0 6.2 41.9 19.8 19.7 8.2 11.5 17.7 GAV 38.7 17.3 13.1 62.7 23.7 25.7 10.8 15.4 25.9 $\widetilde{A}$ 40.2 17.4 13.9 68.2 28.1 27.9 11.2 16.7 28.0 Vanilla 18.5 7.5 5.8 54.3 7.2 11.7 3.4 7.6 14.5 N-Only 19.8 9.4 4.5 47.6 21.8 22.1 9.3 12.6 18.4 GAV 39.1 17.2 12.0 68.6 26.4 28.9 12.2 17.4 27.7 $\widetilde{A}$ 39.0 16.6 11.7 72.5 30.9 30.6 12.4 18.3 29.0
5.3.2 Impact of Architectural Reduction.
@@
Table 4: Investigating architectural reduction’s impact. We compare CLIP’s default encoder architecture, denoted as Vanilla, with a modified setting described in Sec. 4.3, termed Reduced, where the self-attention module’s output directly serves as the final encoder block’s output. Throughout this experiment, we utilize CLIP’s default self-attention module. To ensure clarity regarding the effects, we present results both with and without mask refinement.
Architecture PAMR V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg Vanilla 18.6 7.8 6.5 49.1 6.7 11.2 3.2 7.2 13.8 Reduced 37.5 22.3 23.2 81.4 20.2 24.9 11.6 18.9 30.0 Vanilla 18.5 7.5 5.8 54.3 7.2 11.7 3.4 7.6 14.5 Reduced 37.6 22.5 23.3 81.7 20.3 25.1 11.6 19.2 30.2

### v2

@@ 3.2 Limitations of CLIP for image segmentation
As previously stated, localization holds a pivotal role in segmentation tasks, yet it is often overlooked in ViTs [12]. As described in [12], unlike CNNs, most of the operations in ViTs are global and the locality of patches is mostly not taken into account. Within ViT, and consequently, in CLIP’s visual encoder, positional information is integrated into the network through 1D positional embeddings added to the input at the first block. Recent evidence points to the possible suboptimal performance of 1D embeddings in image data, and thus alternative mechanisms are sought after [25, 9]. Furthermore, positional embeddings utilization is restricted only to the first encoder block, potentially diminishing their relevance in subsequent layers. This oversight could be detrimental in dense prediction tasks, where patches contain important contextual information, and such information remains valuable throughout the network’s depth. Please note that simply incorporating the pre-trained positional embeddings at all the layers is not feasible in the training-free scenario since such additions would change the distribution of the data that each layer expects and requires fine-tuning. In light of the arguments presented, we assert that explicitly attending to the neighbourhood of each patch is imperative within the context of segmentation. Moreover, recent studies have noted that the final encoder block in CLIP’s ViT disrupts spatial information, hindering dense prediction tasks [52]. Indeed, CLIP’s visual encoder is trained to emphasize the [CLS] token’s embedding (global embedding), while the outputs at other locations (i.e., embeddings of patches) are not optimally structured for tasks such as semantic segmentation.
@@ 4 Neighbour-Aware CLIP
The pre-training procedure of CLIP [32] encourages its ViT [12] to learn representations tailored to image-level tasks, thereby compromising its efficacy in dense prediction problems. Given the inherent differences between such tasks and segmentation, coupled with the fact that the output of the ViT for patch tokens has not been explicitly trained during CLIP’s pre-training, it struggles to generalize effectively to pixel-wise prediction scenarios. This underscores the necessity for targeted adjustments in the original CLIP model to accommodate the nuances of semantic segmentation. Our study delves into these specific components that potentially hinder CLIP’s segmentation performance and proposes minimal alterations to its overall framework, without changing any of the network’s parameters. The precise modifications are detailed in this section.
@@ 4.1 Introducing spatial consistency
In Sec. 3.2, we underscored the importance of explicitly attending to the locality of each patch and highlighted the inadequacy of vanilla ViT’s positional embeddings in achieving this. In this section, we introduce our simple method for enforcing explicit spatial attention to each patch’s neighbourhood. In particular, we augment the attention map information with an unnormalized multivariate (in our case, $2$-dimensional) Gaussian kernel, which can be defined as:
@@
Assuming $\boldsymbol{\Sigma}=\sigma^{2}I$, we can rewrite the kernel as
@@
$\begin{gathered}\omega((i,j);\sigma)_{mn}=\phi((m,n);{(i,j)},\sigma),\\ \phantom{,}\forall m\in\{1,2,\dots,h\},\;\forall n\in\{1,2,\dots,w\},\end{gathered}$ (11)
with the maximum value at $(i,j)$, gradually decreasing with distance from $(i,j)$ (see $\omega$ in Fig. 2 for elaboration).
As a corner case, let us assume that the attention map’s logits have been explicitly set to 00 and we set
@@
so that the attention will be purely paid to the neighbourhood of patches. As we show in our empirical validation (termed Neighbourhood Only in Sec. 5.3), by just doing so we can observe a leap in performance compared to CLIP.
Examining Eq. 12, we observe that the patch information is not utilized as $\omega$ replaces $\text{sim}_{ij}$ in Eq. 7, leading to image-independent attention. This prompts us to move beyond the corner case and include the similarity information:
@@ 4.2 Measure of similarity
Transformers [37] are complex deep neural networks and providing a definitive explanation of their inner workings remains challenging. However, from an intuitive standpoint, we can offer the following interpretations for query, key, and value vectors. The query vector denotes what a patch looks for; the key vector signifies what it represents; and the value vector displays what it has to offer. Guided by these descriptions, we deviate from the standard similarity measure ($\mathbf{q}\mathbf{k}^{\top}$) in our self-attention module. This deviation is motivated by the discrepancy between what we want the model to look for (i.e., accurate patch-level predictions) and what it had been trained to look for during pre-training. In semantic segmentation, we need to focus on the nature of each patch and this naturally leads us to shift our attention toward using the key vectors. Thus, we opted to use $\mathbf{k}\mathbf{k}^{\top}$ scores in our similarity measure instead, resulting in $\text{sim}_{ij}={\mathbf{k}\mathbf{k}_{ij}}/{\sqrt{d}}$. By doing so, patches that represent similar information (as portrayed by their key vectors) attend to each other’s value vectors with high intensity. This simple change in perspective brings considerable performance improvement compared to CLIP, as shown in our empirical validation (termed Key-Key Similarity in Sec. 5.3).
Considering the changes proposed up until now, we rewrite Eq. 13 as (see Fig. 2 for more insight):
$\phantom{.}\widetilde{A}_{ij}=\mathrm{softmax}\left(\frac{\mathbf{k}\mathbf{k}_{ij}}{\sqrt{d}}+\omega((i,j);\sigma)\right)\mathbf{v}.$ (15)
[…]
4.3 Eliminating image-level specialized units
@@
in which $L$ denotes the index of the last encoder block. We refer to this structure as the Reduced architecture.
@@ 5 Experiments
Figure 3: Qualitative results (segmentation maps) on PASCAL Context (59) [29] for CLIP [32], SCLIP [38], and our method.
Table 1: Quantitative evaluation on 6 datasets and 2 of their variants. The first 3 benchmarks (V21, PC60, and C-Obj) contain a background category, while the subsequent ones do not. The Fair column indicates whether a method’s comparison to ours is equitable, i.e., conducted without leveraging additional knowledge. The Post. column shows whether an approach contains a post-processing step for mask refinement. For elucidation on abbreviated benchmark names, please refer to Sec. 5.1.
Method Fair Post. V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg
CLIP [32] ICML’21 18.6 7.8 6.5 49.1 6.7 11.2 3.2 5.7 13.6
@@
GroupViT [42] CVPR’22 52.3 18.7 27.5 79.7 18.5 23.4 10.4 15.3 30.7
CLIP Surgery [21] arXiv’23 41.2 30.5 - - 31.4 - 12.9 21.9 -
CLIP-DIY [41] WACV’24 59.0 - 30.4 - - - - - -
GEM [4] CVPR’24 46.2 - - - - 32.6 15.7 - -
SCLIP [38] ECCV’24 59.1 30.4 30.5 80.4 32.2 34.2 16.1 22.4 38.2
NACLIP Ours 58.9 32.2 33.2 79.7 35.5 35.2 17.4 23.3 39.4
@@
FreeSeg-Diff [11] arXiv’24 53.3 - 31.0 - - - - - -
@@
PnP-OVSS [27] CVPR’24 51.3 - 36.2 - - 28.0 14.2 17.9 -
SCLIP [38] ECCV’24 61.7 31.5 32.1 83.5 34.1 36.1 17.8 23.9 40.1
NACLIP Ours 64.1 35.0 36.2 83.0 38.3 38.4 19.1 25.7 42.5
@@
NACLIP 64.1 38.4 51.3 54.8 34.9 44.9 57.9 36.4 47.2
5.1 Experimental setup
Datasets.
We evaluate our method on the following segmentation benchmarks, whose names are abbreviated (in parentheses) to conserve table space: PASCAL VOC 2012 (V21) [13], ADE20K-150 (ADE) [51], PASCAL Context (PC60) [29], COCO-Stuff (C-Stf) [5], Cityscapes (City) [10], COCO-Object (C-Obj) [23]. Additionally, alongside the original benchmarks on these datasets, we follow [38] and evaluate on variants of PASCAL VOC 2012 (V20) and PASCAL Context (PC59) in which the background class is removed from the evaluation. Furthermore, input images are resized to have a shorter side of 336 (560 for Cityscapes [10], because of high-resolution images), and following prior works [38, 3, 21, 7, 42], a slide inference is carried out with a $224\times 224$ window and a stride of 112.
Baselines.
We compare our method to a set of relevant works in OVSS, including: MaskCLIP [52], ReCo [33], CLIP Surgery [21], SCLIP [38], GEM [4], CLIP-DIY [41], FreeSeg-Diff [11], FOSSIL [3], and PnP-OVSS [27]. We also include a few influential weakly supervised OVSS methods, such as GroupViT [42] and TCL [7] in our comparison. Furthermore, since vanilla CLIP [32] can be adapted for semantic segmentation, we include it as a baseline in our comparison tables as well.
Implementation details.
In our experiments, we employ pre-trained CLIP-ViT [32]. Unless mentioned otherwise, we use the ViT-B/16 backbone ($16\times 16$ patch size) comprising 12 visual encoder blocks. Since our approach operates in a training-free manner, we exclusively utilize the frozen CLIP model without any optimization. The standard deviation of the Gaussian kernel, $\sigma$ in Eq. 10, is set to 5, a choice further elaborated upon in Appendix A. OVSS methods often incorporate a mask refinement step [3, 38, 7, 33, 27, 11], such as DenseCRF [20] or pixel-adaptive mask refinement (PAMR) [1]. For our method, we opt for PAMR as it is lighter and more efficient. We also report results without this refinement step. We resort to mIoU for the evaluation metric across all experiments.
5.2 Main results
Comparison to training-free OVSS methods.
In Tab. 1, we evaluate our proposed method against baselines mentioned in Sec. 5.1. Although most methods are presented as training-free OVSS, it is important to note that they vary in underlying characteristics. Specifically, within the subset of methods that refrain from fine-tuning, some leverage auxiliary pre-trained models [33, 3, 41, 27, 11]. Given our primary focus on achieving training-free OVSS exclusively through a frozen CLIP model, comparing our method to those utilizing additional weights and data, while informative, may not be entirely equitable. Hence, we explicitly denote fair comparisons in Tab. 1. Analyzing the results, our approach outperforms state-of-the-art training-free OVSS methods in 7 out of 8 benchmarks. This signifies the effectiveness of our proposed architecture and formulation of CLIP’s visual encoder. The improvement trend is also noticeable among methods that opted not to perform mask refinement steps compared to our method without post-processing. Note that results for newer methods [3, 41, 4, 21, 27, 11] are extracted from their respective manuscripts, hence might not include experiments across all benchmarks, whereas other results have been derived from [38].
Robustness to visual backbones.
Table 2 reports the segmentation results of NACLIP using different CLIP-ViT backbones, namely ViT-B/16 (default), ViT-B/32, and ViT-L/14. Compared to concurrent approaches [38, 4], our method is more robust to the backbone choice. For example, SCLIP [38] experiences a significant performance drop of over 12% on average when using ViT-L/14 instead of the default backbone, whereas the degradation is about 4% in our case. Moreover, our method outperforms GEM [4] by over 10% across all settings, making NACLIP a more robust solution compared to [38, 4].
Visual examples.
Figure 3 presents several segmentation maps generated by our method on the PASCAL Context (59) dataset [29]. These visualizations demonstrate NACLIP’s significant performance enhancement over CLIP. Furthermore, SCLIP seems to find difficulties in identifying the limits of objects properly, sometimes confusing nearby concepts (e.g., the first row). Besides, while SCLIP tends to focus mainly on similar patches without explicitly considering the surrounding context, our method maintains a local contextual understanding. This is evident in the second row (the bus image) where SCLIP misclassifies trees, likely due to its limited attention to nearby objects such as the road. More examples are available in the Appendix B.
@@ 5.3 Ablation study
Effect of altering the self-attention module.
In Sec. 4.1, we discussed the rationale behind introducing spatial consistency to attention maps and described our approach. Besides, we defined the Neighbourhood Only (N-Only in the table) setting, where patches exclusively attend to their neighbours, disregarding similarities. Furthermore, we defined the Key-Key Similarity (KK-Sim in the table) setting in Sec. 4.2, in which $\mathbf{k}\mathbf{k}^{\top}$ is used as the similarity measure. Results from the comparative analysis (Tab. 3) between N-Only and Vanilla, as well as between KK-Sim and Vanilla underscore the efficacy of the modifications proposed in Secs. 4.1 and 4.2, respectively.
Table 3: Ablation on the effect of spatial consistency and similarity measure. Vanilla refers to CLIP’s self-attention module, $\widetilde{A}$ has been defined in Eq. 15, and the rest of the settings have been mentioned in Sec. 5.3. Throughout this experiment, we adhere to the default CLIP encoder architecture, retaining all architectural elements of the encoder block.
Atten. V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg
Vanilla 18.6 7.8 6.5 49.1 6.7 11.2 3.2 5.7 13.6
N-Only 38.0 17.0 16.1 65.9 21.4 23.9 10.2 14.6 25.9
KK-Sim 36.0 15.6 10.9 64.9 24.8 26.2 9.6 15.1 25.4
$\widetilde{A}$ 40.2 17.4 13.9 68.2 28.1 27.9 11.2 16.5 27.9
Impact of architectural reduction.
@@
Table 4: Investigating architectural reduction’s impact. We compare CLIP’s default encoder architecture, denoted as Vanilla, with the Reduced setting described in Sec. 4.3, where the self-attention module’s output directly serves as the final encoder block’s output. Throughout this experiment, we utilize CLIP’s default self-attention module.
Arch. V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg
Vanilla 18.6 7.8 6.5 49.1 6.7 11.2 3.2 5.7 13.6
Reduced 37.5 22.3 23.2 81.4 20.2 24.9 11.6 16.7 29.7